Calculate how much (as grams) the following elements are needed to produce a charge equal to the Faraday’s constant. Electron charge is 1.602*10^–19 As
a) Li (using Li/Li+)
b) Zn (using Zn/Zn2+)
1
Expert's answer
2012-04-19T07:38:17-0400
m = (Q/F)*(M/z)

m
is the mass of the substance liberated at an electrode in grams
Q is the total electric charge passed through the substance
F = 96,485 C mol−1 is the Faraday constant
M is the molar mass of the substance
z is the valency number of ions of the substance (electrons transferred per ion)

For Li

m = 1.602*10–19/96,485 * 6.94/1 = 1.15*10–20

for Zn

m = 1.602*10–19/96,485 * 65.38/2 = 5.4277*10-20